Solve equation:
(x+1/x-1) + (2/x)= (x/x+1)
How do I solve this?

Answer by stanbon(48558)

Solve equation:
(x+1/x-1) + (2/x)= (x/x+1)
-------------------------------
Multiply thru by x(x-1)(x+1) to get:
x(x+1)^2 + 2(x-1)(x+1) = x^2(x-1)
x(x^2 + 2x + 1) + 2(x^2-1) = x^3-x^2
x^3 + 2x^2 + x + 2x^2-2 = x^3-x^2

5x^2 + x -2 = 0
x = [-1 +- sqrt(1 - 4*5*-2)]/10
x = [-1 +- sqrt(41)]/10
==========================================
Cheers,
Stan H.